# Control-Aware Trajectory Predictions for Communication-Efficient Drone Swarm Coordination in Cluttered Environments

### v1

@@ Abstract
Swarms of Unmanned Aerial Vehicles (UAV) have demonstrated enormous potential in many industrial and commercial applications. However, before deploying UAVs in the real world, it is essential to ensure they can operate safely in complex environments, especially with limited communication capabilities. To address this challenge, we propose a control-aware learning-based trajectory prediction algorithm that can enable communication-efficient UAV swarm control in a cluttered environment. Specifically, our proposed algorithm can enable each UAV to predict the planned trajectories of its neighbors in scenarios with various levels of communication capabilities. The predicted planned trajectories will serve as input to a distributed model predictive control (DMPC) approach. The proposed algorithm combines (1) a trajectory compression and reconstruction model based on Variational Auto-Encoder, (2) a trajectory prediction model based on EvolveGCN, a graph convolutional network (GCN) that can handle dynamic graphs, and (3) a KKT-informed training approach that applies the Karush–Kuhn–Tucker (KKT) conditions in the training process to encode DMPC information into the trained neural network. We evaluate our proposed algorithm in a funnel-like environment. Results show that the proposed algorithm outperforms state-of-the-art benchmarks, providing close-to-optimal control performance and robustness to limited communication capabilities and measurement noises.
@@ I Introduction
Recently, several works have attempted to develop communication-efficient control strategies for UAV swarms and other decentralized multi-robot systems, aiming to reduce communication costs for exchanging state information crucial for decision-making and control [8, 12, 13, 21]. Existing research can be divided into two categories. The first category predicts planned trajectories based on historical states measured by onboard sensors [34, 36, 38, 35, 33]. For example, [8] developed a learning-based approach based on recurrent neural networks to predict the planned trajectories of each UAV, which does not require communications. However, it is assumed that each UAV can sense the real-time states of all other UAVs, which may not be possible in cluttered environments. The second category seeks to compress the communicated messages on planned trajectories to reduce the bandwidth requirement [37, 39, 40]. For example, [12] proposed a UAV path discretization and compression algorithm that simultaneously optimizes the selection of waypoints and the subpath trajectories while ensuring the optimal performance of trajectory optimization. However, the proposed algorithm only focused on each individual path without exploiting their correlations. Reference [13] developed a neural network-based mean-field approximation approach to approximate the solution to the Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck-Kolmogorov (FPK) equations, which can effectively minimize communication. To avoid violating required conditions, this work introduced Federated Learning to enable the sharing of model parameters between UAVs. However, this work requires the online training and sharing of the neural network parameters, which even at low communication frequency, can be a large amount of data to exchange.
@@
From the literature mentioned above, we have identified several research gaps in existing research on communication-efficient UAV swarm coordination. First, existing works either attempt to predict trajectories from historical states or compress communicated messages on planned trajectories but rarely combine both approaches such that these two types of information can compensate for each other. Second, many existing works have made strong assumptions of sensing range or reliable communications and have yet to test their proposed algorithms in scenarios with cluttered environments. Third, existing works rarely consider information about the decision-making process of the UAV swarm, e.g., optimization models, which can be common knowledge for all UAVs in the same swarm. Integrating such knowledge can help ensure that the predicted trajectories are more physically meaningful.
@@
In this paper, we address the aforementioned research gaps by developing a learning-based algorithm that enables each UAV to predict the trajectories of its neighbors in cluttered environments with various levels of communication capabilities. Our algorithm combines three types of information: (i) the historical states of each UAV and its neighbors, (ii) compressed messages communicated from neighbors, if available, and (iii) the information about the decision-making process of each UAV, characterized by a convex optimization-based DMPC approach. The main part of our algorithm combines the first two types of information by integrating two models in a Bayesian framework: a Variational Auto-Encoder (VAE)-based trajectory compression and reconstruction model and a graph convolutional network (GCN)-based trajectory prediction model. Our trajectory prediction model can handle the dynamics scenarios with time-varying swarm alignment, where the neighbors of a given UAV constantly change. This is achieved by leveraging an EvolveGCN architecture [7] that can track the change in the underlying graph. The two models are trained together via a Karush–Kuhn–Tucker (KKT) conditions-informed method [10] to encode the information of the convex optimization model upon which each UAV makes decisions (i.e., the third type of information).
@@
The main contributions of this paper are two-fold. First, we devise a learning-based algorithm to predict the trajectories of surrounding UAVs in cluttered environments with various levels of communication capabilities (bandwidths, packet loss rates, lower frequencies, etc.), which is achieved by combining an EvolveGCN-based trajectory prediction model with a VAE-based trajectory compression and reconstruction model in a Bayesian framework. Second, we train the developed neural network via a KKT-informed approach that enables the trajectory prediction algorithm to encode the information about the decision-making process of UAVs to improve control performance. To the best of our knowledge, no existing trajectory prediction algorithms have explicitly encoded knowledge about the decision-making processes of the agents.
@@ II Problem Statement
As shown in Figure 1, we consider a swarm of UAVs indexed by $i\in\mathcal{V}$ in a funnel-like environment where the obstacles, indexed by $m\in\mathcal{M}$, are convex and modeled as 3-dimensional ellipsoids, each parameterized by a center point $\bm{C}_{m}\in\mathbb{R}^{3}$ and an affine transformation matrix $\bm{E}_{m}\in\mathbb{R}^{3\times 3}$. The considered time horizon is discretized as a set of intervals $\mathcal{K}=\{1,2,\cdots,K\}$ of a given size $\Delta T$ indexed by $k$. The state of UAV $i\in\mathcal{V}=\{1,2,\cdots,n\}$ at time step $k\in\mathcal{K}$ is represented by $\bm{x}_{i}(k)=\left(\bm{p}_{i}(k),\bm{v}_{i}(k)\right)\in\mathbb{R}^{6}$, where $\bm{p}_{i}(k)\in\mathbb{R}^{3}$ and $\bm{v}_{i}(k)\in\mathbb{R}^{3}$ denote its location and velocity at time step $k$, respectively. The dynamics of UAV $i$ can be modeled as a discrete linear system with an underlying position controller:
@@
Each UAV $i$ can collect two types of information: (1) the real-time states of other UAVs within a sensing range, denoted by $\bm{x}_{j}(t)$ for the states of UAV $j$ at time step $t$, and (2) the messages sent by other UAVs within a communication range, denoted by $\bm{z}_{j}(t)$ for the message sent by UAV $j$ at time step $t$. For brevity, we assume that the communication graph and the sensing graph are identical, both represented by $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where edge $e=(i,j)\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ indicates that UAV $j$ is within the sensing/communication range of UAV $i$. Such an assumption is made without loss of generality since the developed algorithm can be extended to distinguish communication and sensing graphs. We further make the following remarks regarding the sensing/communication graph $\mathcal{G}$. First, we consider such a graph to be dynamic, specifically with a time-invariant node set $\mathcal{V}$ (i.e., fixed swarm) and a time-varying edge set $\mathcal{E}$ (i.e., dynamic neighbors). Such an assumption is realistic since the alignment of UAVs can change dynamically, especially in cluttered environments. Second, the messages sent by UAV $j$ to UAV $i$ can contain possibly compressed information about UAV $j$’s planned trajectories, which will be used to facilitate the decision-making of UAV $i$. Note that the trajectory information is compressed to cater to the limited bandwidth of the communication channel. To simplify the implementation, we assume that each UAV sends identical structured messages to all its neighbors on the communication graph. Nevertheless, this assumption can be easily extended such that UAVs can share customized messages with their neighbors under specific conditions.
@@
To better formulate collision-avoidance constraints, the planned trajectory of UAV $i$, $\bm{U}_{i}^{k}$, is parameterized by $l$ 3-dimensional Bezier curves of order $d$, whereby each of these Bezier curves can be uniquely characterized by a set of $d+1$ 3-dimensional control points. Consequently, the planned trajectory $\bm{U}_{i}^{k}$ is characterized by a vector $\bm{w}_{i}^{k}\in\mathbb{R}^{3l(d+1)}$ formed by $l(d+1)$ 3-dimensional control points. The relation between the planned trajectory and the control points can be described as $\bm{U}_{i}^{k}=\bm{F}_{i}\bm{w}_{i}^{k}$, where $\bm{F}_{i}$ indicates the coefficients of the Bezier curves.
With the aforementioned trajectory parametrization, the embedded QP problem of the DMPC scheme for UAV $i$ can be described in (2)-(8). For brevity, here we provide the general forms and insights of the QP formulation. For the specific derivation of the cost terms and constraint matrices, please refer to [28].
$\min_{\bm{w}_{i}^{k},\bm{\epsilon}_{i}^{k},\bm{\delta}_{i}^{k},\bm{\zeta}_{i}^{k}}$ $\leavevmode\nobreak\ J^{k}_{i}\Big{(}\bm{w}_{i}^{k},\bm{\epsilon}_{i}^{k},\bm{\delta}_{i}^{k},\bm{\zeta}_{i}^{k}\Big{)}$ (2)
s.t. ${\bm{A}_{\text{dyn},i}}\bm{w}^{k}_{i}\leq\bm{b}_{\text{dyn},i}$ (3)
$\bm{A}_{\text{cont},i}\bm{w}^{k}_{i}=\bm{b}_{\text{cont},i}$ (4)
@@
$\quad\quad\leq\bm{b}^{k}_{\text{saf-agent},i}\left(\tilde{\bm{U}}_{-i}^{k}\right)$ (5)
${\bm{A}^{k}_{\text{coh},i}}\left(\tilde{\bm{U}}_{-i}^{k}\right)\left[(\bm{w}^{k}_{i})^{T},(\bm{\delta}^{k}_{i})^{T}\right]^{T}\leq\bm{b}^{k}_{\text{coh},i}\left(\tilde{\bm{U}}_{-i}^{k}\right)$ (6)
${\bm{A}^{k}_{\text{saf-obs},i}}[({\bm{w}}^{k}_{i})^{T},(\bm{\zeta}^{k}_{i})^{T}]^{T}\leq\bm{b}^{k}_{\text{saf-obs},i}$ (7)
$\bm{\epsilon}_{i}^{k},\bm{\delta}_{i}^{k},\bm{\zeta}_{i}^{k}\geq 0$ (8)
where the objective function is a quadratic function involving (i) migration cost that penalizes the distance to a pre-defined migration point, (ii) inter-agent and agent-obstacle collision costs penalizing the UAV of getting too close to its neighbors and obstacles, (iii) cohesion cost to ensure that adjacent UAVs stay closer than a threshold cohesion distance, and (iv) control effort cost to minimize the energy spent to execute the control action. Constraint (3) ensures the feasibility of the dynamics given by the planned trajectory parameterized as the Bezier curve. Constraint (4) imposes $C^{2}$-continuity requirement on the derived Bezier curve. Constraints (5) and (6), respectively, indicate the collision avoidance and cohesion between UAV $i$ and its surrounding UAVs to bound the distance between them. Constraint (7) indicates the collision avoidance between UAV $i$ and obstacles. Note that constraints (5)-(7) are linearized from nonlinear constraints and relaxed with relaxation terms $\bm{\epsilon}^{k}_{i}$, $\bm{\delta}^{k}_{i}$, and $\bm{\zeta}^{k}_{i}$ minimized in the objective function. Constraint (8) requires these relaxation terms to be nonnegative.
Further note that constraints (5) and (6) depend on the predicted trajectories of the surrounding UAVs, i.e., $\tilde{\bm{U}}_{-i}^{k}$. Hence, the quality of decisions, especially on inter-agent safety and cohesion, heavily depends on the accuracy of such a prediction. Originally in [28], such predictions are assumed to be planned trajectories always available with moderate noises, as they can be obtained through communicating with surrounding UAVs at each time step. However, since the communication channel may have limited bandwidth and frequency, as well as possibility of packet loss, it may not be feasible to share the entire trajectory at each time step. Therefore, it is important to develop a communication-efficient and robust method for trajectory prediction, which will be developed in Section III.

### v2

@@ Abstract
Swarms of Unmanned Aerial Vehicles (UAV) have demonstrated enormous potential in many industrial and commercial applications. However, before deploying UAVs in the real world, it is essential to ensure they can operate safely in complex environments, especially with limited communication capabilities. To address this challenge, we propose a control-aware learning-based trajectory prediction algorithm that can enable communication-efficient UAV swarm control in a cluttered environment. Specifically, our proposed algorithm can enable each UAV to predict the planned trajectories of its neighbors in scenarios with various levels of communication capabilities. The predicted planned trajectories will serve as input to a distributed model predictive control (DMPC) approach. The proposed algorithm combines (1) a trajectory prediction model based on EvolveGCN, a Graph Convolutional Network (GCN) that can handle dynamic graphs, which is further enhanced by compressed messages from adjacent UAVs, and (2) a KKT-informed training approach that applies the Karush-Kuhn-Tucker (KKT) conditions in the training process to encode DMPC information into the trained neural network. We evaluate our proposed algorithm in a funnel-like environment. Results show that the proposed algorithm outperforms state-of-the-art benchmarks, providing close-to-optimal control performance and robustness to limited communication capabilities and measurement noises.
@@ I Introduction
Recently, several works have attempted to develop communication-efficient control strategies for UAV swarms and other decentralized multi-robot systems, aiming to reduce communication costs for exchanging state information crucial for decision-making and control [21, 22, 23, 24, 25]. Existing research can be divided into two categories. The first category predicts planned trajectories based on historical states measured by onboard sensors [26, 27, 28, 29, 30]. For example, Ref. [21] developed a learning-based approach based on recurrent neural networks to predict the planned trajectories of each UAV, which does not require communications. However, this method overlooks the frequently changing topology caused by the complex interactions among UAVs in cluttered environments. The second category seeks to compress the communicated messages on planned trajectories to reduce the bandwidth requirement [31, 32, 33]. For example, Ref. [22] proposed a UAV path discretization and compression algorithm that simultaneously optimizes the selection of waypoints and the subpath trajectories while ensuring the optimal performance of trajectory optimization. However, the proposed algorithm only focused on each individual path without exploiting their correlations. Ref. [23] developed a neural network-based mean-field approximation approach to approximate the solution to the Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck-Kolmogorov (FPK) equations, which can effectively minimize communication. To avoid violating required conditions, this work introduced Federated Learning to enable the sharing of model parameters between UAVs. However, this work requires the online training and sharing of the neural network parameters, which even at low communication frequency, can be a large amount of data to exchange.
@@
From the literature mentioned above, we have identified two research gaps in existing research on communication-efficient UAV swarm coordination. First, existing works either attempt to predict trajectories from historical states or compress communicated messages on planned trajectories but rarely combine both approaches such that these two types of information can compensate for each other. Moreover, many existing works overlook the changing topology of UAV swarms or have made strong assumptions of reliable communications and have yet to test their proposed algorithms in scenarios with cluttered environments. Second, existing works rarely consider information about the decision-making process of the UAV swarm, e.g., optimization models, which can be common knowledge for all UAVs in the same swarm. Integrating such knowledge can help ensure that the predicted trajectories are more physically meaningful.
@@
In this paper, we address the aforementioned research gaps by developing a learning-based algorithm that enables each UAV to predict the trajectories of its neighbors in cluttered environments with various levels of communication capabilities. Our algorithm combines three types of information: (i) the historical states of each UAV and its neighbors, (ii) compressed messages communicated from neighbors, if available, and (iii) the information about the decision-making process of each UAV, characterized by a convex optimization-based DMPC approach. Our algorithm combines the first two types of information by integrating two modules in a Bayesian framework: a graph convolutional network (GCN)-based trajectory prediction module and a Variational Auto-Encoder (VAE)-based trajectory compression and reconstruction module. Our trajectory prediction module can handle dynamic scenarios with time-varying swarm alignment, where the neighbors of a given UAV constantly change. This is achieved by leveraging an EvolveGCN architecture [34] that can track the change in the underlying graph. The third type of information, i.e., the information of the convex optimization model upon which each UAV makes decisions, is encoded in a Karush–Kuhn–Tucker (KKT) conditions-informed training method [35] that trains the two modules together.
@@
Statement of Contribution. The main contributions of this paper are two-fold. First, we devise a learning-based algorithm to predict the trajectories of surrounding UAVs in cluttered environments with various levels of communication capabilities (bandwidths, packet loss rates, lower frequencies, etc.), which is achieved by combining an EvolveGCN-based trajectory prediction model with a VAE-based trajectory compression and reconstruction model in a Bayesian framework. Second, we train the developed neural network via a KKT-informed approach that enables the trajectory prediction algorithm to encode the information about the decision-making process of UAVs to improve control performance. To the best of our knowledge, no existing trajectory prediction algorithms have explicitly encoded knowledge about the decision-making processes of the agents.
@@ II Problem Statement
As shown in Figure 1, we consider a swarm of UAVs indexed by $i\in\mathcal{V}$ in a funnel-like environment where the obstacles, indexed by $m\in\mathcal{M}$, are convex and modeled as 3-dimensional ellipsoids, each parameterized by a center point $\bm{C}_{m}\in\mathbb{R}^{3}$ and an affine transformation matrix $\bm{E}_{m}\in\mathbb{R}^{3\times 3}$. The goal of the UAV swarm is to safely navigate to a predefined migration point (represented by the red point in Figure 1). The considered time horizon is discretized as a set of intervals $\mathcal{K}=\{1,2,\cdots,K\}$ of a given size $\Delta T$ indexed by $k$. The state of UAV $i\in\mathcal{V}=\{1,2,\cdots,n\}$ at time step $k\in\mathcal{K}$ is represented by $\bm{x}_{i}(k)=\left(\bm{p}_{i}(k),\bm{v}_{i}(k)\right)\in\mathbb{R}^{6}$, where $\bm{p}_{i}(k)\in\mathbb{R}^{3}$ and $\bm{v}_{i}(k)\in\mathbb{R}^{3}$ denote its location and velocity at time step $k$, respectively. The dynamics of UAV $i$ can be modeled as a discrete linear system with an underlying position controller:
@@
Each UAV $i$ can collect two types of information: (1) the real-time states of other UAVs within a sensing range, denoted by $\bm{x}_{j}(t)$ for the states of UAV $j$ at time step $t$, and (2) the messages sent by other UAVs within a communication range, denoted by $\bm{z}_{j}(t)$ for the message sent by UAV $j$ at time step $t$. For sensing, we assume each UAV can fully sense the entire swarm, so the sensing graph $\mathcal{G}_{\text{sen}}=(\mathcal{V}_{\text{sen}},\mathcal{E}_{\text{sen}})$ is a complete graph. This can be achieved through the multi-sensor fusion of LiDAR, cameras, ultra-wideband, etc., which has been demonstrated effective even in cluttered environments [36, 37]. For communication, each UAV is assumed to be able to communicate with the nearest up to four UAVs within a given range, forming a communication graph $\mathcal{G}_{\text{comm}}=(\mathcal{V}_{\text{comm}},\mathcal{E}_{\text{comm}})$. Let neighbors $\mathcal{N}_{i}$ denote the set of UAVs that UAV $i$ can communicate with. We make the following remarks regarding the communication graph $\mathcal{G}_{\text{comm}}$. First, we consider the nearest up to four neighbors because the movement of the UAV is mainly influenced by these immediate neighbors. Such a treatment helps balance the swarm performance, computational complexity of the controller, and communication burden [17]. Second, we consider the communication graph $\mathcal{G}_{\text{comm}}$ to be dynamic, with both the node set $\mathcal{V}_{\text{comm}}$ and edge set $\mathcal{E}_{\text{comm}}$ varying over time to represent changing neighbors. Such an assumption is realistic since the alignment of UAVs can change dynamically, especially in cluttered environments. Third, the messages sent by UAV $j$ to UAV $i$ can contain possibly compressed information about UAV $j$’s planned trajectories, which will be used to facilitate the decision-making of UAV $i$. Note that the trajectory information is compressed to cater to the limited bandwidth of the communication channel. To simplify the implementation, we assume that each UAV sends identical structured messages to all its neighbors on the communication graph. Nevertheless, this assumption can be easily extended such that UAVs can share customized messages with their neighbors under specific conditions.
@@
To better formulate collision-avoidance constraints, the planned trajectory of UAV $i$, $\bm{U}_{i}^{k}$, is parameterized by $l$ 3-dimensional Bézier curves of order $d$, whereby each of these Bézier curves can be uniquely characterized by a set of $d+1$ 3-dimensional control points. Consequently, the planned trajectory $\bm{U}_{i}^{k}$ is characterized by a vector $\bm{w}_{i}^{k}\in\mathbb{R}^{3l(d+1)}$ formed by $l(d+1)$ 3-dimensional control points. This parameterization allows us to define continuous trajectories in a polynomial form $f(\bm{w}_{i}^{k})$. The relation between the planned trajectory and the control points can be described as $\bm{U}_{i}^{k}=\bm{F}_{i}\bm{w}_{i}^{k}$, where $\bm{F}_{i}$ indicates the coefficients of the Bézier curves.
With the aforementioned trajectory parametrization, the embedded QP problem of the DMPC scheme for UAV $i$ can be described in Eq.(2a)-(2g), as inspired by [17].
$\min_{\bm{w}_{i}^{k},\bm{\zeta}_{i}^{k},\bm{\epsilon}_{i}^{k},\bm{\delta}_{i}^{k}}$ $\leavevmode\nobreak\ \sum_{\tau=1}^{P}q_{\text{mig }}\left\|\bm{p}_{i}(k+\tau\mid k)-\bm{p}_{\text{mig }}\right\|_{2}^{2}$
$+l_{\mathrm{saf}}\bm{\zeta}_{im}\left(k_{\mathrm{coll},i}\mid k\right)+q_{\mathrm{saf}}\bm{\zeta}_{im}^{2}\left(k_{\mathrm{coll},i}\mid k\right)$
$+\sum_{j\in\mathcal{N}_{i}}\sum_{\tau=0}^{P-1}\left(l_{\text{saf }}\bm{\epsilon}_{ij}(k+\tau\mid k)+q_{\text{saf }}\bm{\epsilon}_{ij}^{2}(k+\tau\mid k)\right)$
$+\sum_{j\in\mathcal{N}_{i}}\sum_{\tau=0}^{P-1}\left(l_{\text{coh }}\bm{\delta}_{ij}(k+\tau\mid k)+q_{\text{coh }}\bm{\delta}_{ij}^{2}(k+\tau\mid k)\right)$
$+\sum_{\tau=0}^{P-1}q_{\text{eft }}\left\|\frac{d^{2}}{dt^{2}}f(\bm{w}_{i})(k+\tau\mid k)\right\|_{2}^{2}$ (2a)
s.t. ${\bm{A}_{\text{dyn},i}}\bm{w}^{k}_{i}\leq\bm{b}_{\text{dyn},i}$ (2b)
$\bm{A}_{\text{cont},i}\bm{w}^{k}_{i}=\bm{b}_{\text{cont},i}$ (2c)
@@
$\quad\quad\leq\bm{b}^{k}_{\text{saf-agent},i}\left(\tilde{\bm{U}}_{-i}^{k}\right)$ (2d)
${\bm{A}^{k}_{\text{coh},i}}\left(\tilde{\bm{U}}_{-i}^{k}\right)\left[(\bm{w}^{k}_{i})^{T},(\bm{\delta}^{k}_{i})^{T}\right]^{T}\leq\bm{b}^{k}_{\text{coh},i}\left(\tilde{\bm{U}}_{-i}^{k}\right)$ (2e)
${\bm{A}^{k}_{\text{saf-obs},i}}[({\bm{w}}^{k}_{i})^{T},(\bm{\zeta}^{k}_{i})^{T}]^{T}\leq\bm{b}^{k}_{\text{saf-obs},i}$ (2f)
$\bm{\epsilon}_{i}^{k},\bm{\delta}_{i}^{k},\bm{\zeta}_{i}^{k}\geq 0$ (2g)
where $\bm{p}_{\text{mig }}$ is the pre-defined goal position that UAV swarm aims to reach, $k_{\mathrm{coll},i}$ is the time that UAV $i$ would first collide with obstacle $m$ in its predicted horizon if the control actions remain the same. Slack variables $\bm{\zeta},\bm{\epsilon},\bm{\delta}$ represent the violation of the safe agent-obstacle distance, safe inter-agent distance, and desired inter-agent cohesion distance, respectively. The objective function is a quadratic function involving (i) migration cost that penalizes the distance to a pre-defined migration point, (ii) agent-obstacle and inter-agent collision costs penalizing the UAV of getting too close to obstacles and its neighbors, (iii) cohesion cost to ensure that adjacent UAVs stay closer than a threshold cohesion distance, and (iv) control effort cost, characterized by the second derivative of the polynomial Bézier curve (i.e., the acceleration), which is minimized to reduce the energy spent on executing the control action. Constraint (2b) ensures the feasibility of the dynamics given by the planned trajectory parameterized as the Bézier curve. Constraint (2c) imposes $C^{2}$-continuity requirement on the derived Bézier curve. Constraints (2d) and (2e), respectively, indicate the collision avoidance and cohesion between UAV $i$ and its surrounding UAVs to bound the distance between them. Constraint (2f) indicates the collision avoidance between UAV $i$ and obstacles. Note that constraints (2d)-(2f) are linearized from nonlinear constraints using first-order Taylor expansion. The constraint (2g) requires these relaxation terms to be nonnegative.
Further note that constraints (2d) and (2e) depend on the predicted trajectories of the surrounding UAVs, i.e., $\tilde{\bm{U}}_{-i}^{k}$. Hence, the quality of decisions, especially on inter-agent safety and cohesion, heavily depends on the accuracy of such a prediction. Originally in [17], such predictions are assumed to be planned trajectories always available with moderate noises, as they can be obtained through communicating with surrounding UAVs at each time step. However, since the communication channel may have limited bandwidth and frequency, as well as possibility of packet loss, it may not be feasible to share the entire trajectory at each time step. Therefore, it is important to develop a communication-efficient and robust method for trajectory prediction, which will be developed in Section III.